OpenStudy (anonymous):

Identify the roots and state the multiplicity x^6 - 12x^4 + 48x^2 -64 = 0

OpenStudy (anonymous):

i put it in the calculator and I see that the zeros are at 2 and -2 i am not clear on the multiplicity of this @Chlorophyll

OpenStudy (anonymous):

hi @vishal kothari can you help me understand this?

OpenStudy (anonymous):

@saifoo.khan

OpenStudy (saifoo.khan):

Sorry i dont really know how to do it. But what i think is, we have to suppose the values for let's say x^2 x^2 =y Then equation will look something like, y^3 - 12 y^2 + 48y - 64 = 0

OpenStudy (saifoo.khan):

i think i solved it. lol Now factor this out, y^3 - 12 y^2 + 48y - 64 = 0 (y-4)(y-4)(y-4) = 0 <--- used a calc here. you have y = 4. Now use this value here; x^2 =y

OpenStudy (saifoo.khan):

@JuanitaM , does it make any sense?

OpenStudy (anonymous):

well oddly enough the zeroes are 2 and -2 not 4. thanks anyway

OpenStudy (saifoo.khan):

Yes! right.

OpenStudy (kinggeorge):

What saifoo is saying is that \(y=x^2=4\) thus, \(x=\pm 2\) . Then, since you have degree 6, and each root is either \(\pm 2\), both roots have multiplicity 3.

OpenStudy (saifoo.khan):

i dont get the degree thing.. but i was correct. :D

OpenStudy (anonymous):

yes u were :)

OpenStudy (kinggeorge):

To find the multiplicity more explicitly, take the function\[(x^2-4)(x^2-4)(x^2-4)=0\]Then for each root \(x^2=4\), of which there are three, you have two roots; \(+2, -2\). Since each of these appears three times, each has multiplicity 3.

OpenStudy (saifoo.khan):

Ah, prefect.

OpenStudy (anonymous):

great - but how did you off the back see x^2 -4 we graphed and that is how i knew the zeros was +-2. I misread the answer so each term is amultiplicy of 3 with a total of 6

OpenStudy (kinggeorge):

When saifoo wrote \((y-4)(y-4)(y-4)\), this is equal to \((x^2 -4)(x^2 -4)(x^2 -4)\) since \(y=x^2\). That's where I got the \((x^2 -4)(x^2 -4)(x^2 -4)\) from.

OpenStudy (anonymous):

i think it is thee cube of first and last terms -